Outils pour utilisateurs

Outils du site


new_measurement_cern_0318

Mesure aux cern

telnet 128.141.143.237 1527

Measurement with standard Gas

Day run start end HVapp I bot Itop Thr T[C] P[hPa] H[%]
1 Mar739327 12:– 14:18 7000 0.49 0.83 500 22.6 945 18
1 Mar739331 14:33 16:48 6900 0.43 0.77 500 22.7 946 16
1 Mar739332 16:54 7:01 6800 0.36 0.77 500 22.5 944 16
1,2 Mar739333 7:04 06:55 5000 0.04 0.06 500 22.15 943 16
2 Mar739334 7:01 10:16 6700 0.47 0.65 500 22.3 939 13
2 Mar739335 10:20 14:23 6600 0.31 0.54 500 22.3 942 16
2 Mar739338 14:25 16:44 6500 0.31 0.50 500 22.1 945 16
2 Mar739340 16:52 19:06 6200 0.23 0.41 500 22.1 945 16
2,3 Mar739342(41) 19:15 09:09 71000.95 1.34 50022.4 94516
3 Mar739343 10:30 Crashed (15:32) (Deleted) 71000.75 1.13 50022.4 94216
3 Mar739345 15:42 Stopped memory leak (Deleted) 71000.96 1.21 50022.4 94616
3 Mar739346 16:19 –:– 71000.96 1.21 50022.4 94616
4 Mar739371 09:45 14:29 (Deleted) 71000.84 1.73 500 22.294318
4 Mar739388 14:3917:29 72001.09 1.63 500 22.794018
4 Mar739395 17:31 19:59 73000.76 1.45 500 22.093818
4 Mar739396 20:05 22:25 74000.56 1.04 500 22.193918

* RUN after gas stop* *

Day run start end HVapp I bot Itop Thr T[C] P[hPa] H[%]
21 Mar739461 10:58 –:– 70000.380.6050022.496618
21 Mar739462 –:– – :– 68000.340.4450022.796818
21 Mar739463 14:30 – :– 68000.380.4822.996616
21 Mar739464 –:– – :– 68000.380.5623.096716
21 Mar739465 –:– – :– 68000.320.5222.196716
21 Mar739466 –:– – :– 68000.340.4722.096816

From run 739346 new config: config_FE1_sdhcall1_5 (no root writing)

RUN: 739342: switched off HV in 2am.

DAQ user 2018/03/02 06:55

I stop run 739333. The noise is reasonable

  • 3 / 2.5 ms on TDC8
  • 5 / 2.5 ms on TDC5

⇒ The noise seen @ higher HV is coming from the chamber

I restart with 6700 V

DAQ user 2018/03/03 15:45

There was acrash of the event builder , most probably due to rootwriter Monday I will supress this root file writing

In the mean time in case of problem:

daqcontrol --jc-kill

daqcontrol --jc-start;daqcontrol --daq-create;daqcontrol --daq-disc;daqcontrol --daq-serv;daqcontrol --daq-init;daqcontrol --daq-conf

/home/acqilc/prepare_cosmic

daqcontrol --daq-start
daqcontrol --trig-resume

Measurement with New Gas System

Scan in threshold

HV applied = 6800 V , Scintillator on TDC 8

P from 966 to 969 (462→465) T ~ 22,22.5 C

Run State DB TDC 8 Thr TDC5 Thr Efficiency (TDC8) Efficiency (2D) Multiplicity Multiplicity (2D) P T HV eff
739462FE1_24CH_17 494 499 80.97+- 1.2080.59+- 1.21 7.0 3.8966 22.57090
739463FE1_24CH_18 497 501 81.49+- 2.2180.84+- 2.24 4.8 2.7966 22.57090
739464FE1_24CH_19 500 504 76.92+- 2.3976.28+- 2.41 3.9 2.4968 22.3
739465FE1_24CH_20 503 507 71.77+- 1.7471.62+- 1.75 3.1 2.1 969 22.0
739466FE1_24CH_21 506 510 68.05+- 2.1267.43+- 2.13 2.5 2.0969 21.8
739467FE1_24CH_22 509 513 61.20+- 2.8260.20+- 2.83 2.3 1.9969 22.5
739468FE1_24CH_23 515 519 57.69+- 2.5054.62+- 2.52 2.0 1.8969 22.5
739469FE1_24CH_24 521 525 51.33+- 3.0845.25+- 3.07 1.9 1.8970 22.3
739470FE1_24CH_25 533 537 39.19+- 2.9532.23+- 2.83 1.9 1.8970 22.1
739472FE1_24CH_26 505 510 72.77+- 1.0672.10+- 1.06 2.7 1.9969 22.2

DAQ user 2018/03/22 05:44

I keep the same setting but rise the HV to 7 kV and make few points

Run State DB TDC 8 Thr TDC5 Thr Efficiency (TDC8) Efficiency (2D) Multiplicity Multiplicity (2D) P T
739473FE1_24CH_26 505 510 92.17+- 1.2791.95+- 1.29 4.3 2.8968 22.4
739474FE1_24CH_20 503 507 93.14+- 1.4592.48+- 1.51 4.6 2.8968 22.4
739475FE1_24CH_19 500 504 94.08+- 1.1393.85+- 1.15 5.4 3.2967 22.5
739476FE1_24CH_18 497 501 95.97+- 1.7795.16+- 1.93 6.8 4.3968 22.6
739477FE1_24CH_26 505 510 90.60+- 0.7190.18+- 0.73 4.2 2.7967 22.7

Scan in HV

Run State DB HV applied Efficiency (TDC8) Efficiency (2D) Multiplicity Multiplicity (2D) P T HV eff
739480FE1_24CH_26 6900 90.73+- 1.6790.73+- 1.67 3.8 2.5967 22.77085
739477FE1_24CH_26 7000 90.60+- 0.7190.18+- 0.73 4.2 2.7967 22.77187
739482FE1_24CH_26 6900 88.47+- 0.6788.16+- 0.68 3.6 2.4963 22.57104
739483FE1_24CH_26 6800 81.49+- 1.1380.98+- 1.14 3.0 2.1958 22.87036
739484FE1_24CH_26 6700 65.67+- 1.5565.25+- 1.55 2.7 1.9957 22.86938
739491FE1_24CH_26 6600 37.09+- 1.6935.99+- 1.68 2.7 2.1955 22.86848
739501FE1_24CH_26 6500 10.68+- 1.36 9.71+- 1.30 3.7 2.9953 22.56748
739517FE1_24CH_26 7000 90.83+- 0.8390.58+- 0.84 4.7 3.1951 21.97267
739519FE1_24CH_26 7100 89.45+- 0.4989.20+- 0.49 5.8 3.7950 21.97377
739520FE1_24CH_26 7200 88.12+- 1.3387.10+- 1.38 6.9 4.3946 22.57518
739521FE1_24CH_22 6900 (Deleted) 88.74+- 0.2888.40+- 0.29 3.5 2.5946 22.57205
739528FE1_24CH_22 6900 87.32+- 0.5486.95+- 0.55 3.3 2.3959 22.57128

Scan in HV Middle position

Run State DB HV applied Efficiency (TDC8) Efficiency (2D) Multiplicity Multiplicity (2D) P T HV eff
739529FE1_24CH_26 6900 90.29+- 1.0685.82+- 1.25 3.3 2.8961 22.57116
739531FE1_24CH_26 6800 85.20+- 1.1478.52+- 1.32 2.7 2.3961 22.07003

DAQ user 2018/03/26 17:26 I patched the code to write only 0.1 % of noise event and only trigger events

739534FE1_24CH_26 6800 83.77+- 1.0877.98+- 1.22 2.7 2.4962 22.17001
739535FE1_24CH_26 6700 66.39+- 1.5258.90+- 1.59 2.1 2.0963 22.36894
739536FE1_24CH_26 6600 34.45+- 1.0028.82+- 0.96 2.2 2.0964 22.46788
739537FE1_24CH_26 7000 90.15+- 0.9887.47+- 1.08 3.7 3.2964 21.87188
739539FE1_24CH_26 7100 82.30+- 1.3279.10+- 1.40 4.1 3.4964 22.17296
739541FE1_24CH_26 7100 85.82+- 1.5083.06+- 1.61 4.1 3.4963 21.97299
739543FE1_24CH_26 6900 XXXX+- 1.5083.06+- 1.61 4.1 3.4963 21.97093

DAQ user 2018/03/28 14:58

Run were stopped to investigate current increased and noise. Moreover on 2 strips (79 & 81)of TDC8 I now see a jump of 2.5 ns to be understood.

The jump is due to the abscence of global reset of all TDC counter so from one reset to the other the TDC channel may be displaced by one clock. Following Imad idea I align all the strips to one and repeat runs in same conditions (6900 V) to check the stability

739560FE1_24CH_26690091.76+- 0.8788.58+- 1.00 3.4 2.995421.87144
739573FE1_24CH_26690090.86+- 2.0586.80+- 2.41 3.5 2.995521.57132
739581FE1_24CH_266900XXXXX86.80+- 2.41 3.5 2.995521.57132

Scan with Large scintillators

Session APRIL1: 1/4

Iuri prepared large scintillators that we put in 1/4 position. It s between the High Radius and central one to guarantee the full acceptance. All 4 Febs are on. We start on 16/04.

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739797FE1_24CH_29710016/04 96121.87306
739800FE1_24CH_29690016/04 96321.67085
739801FE1_24CH_29670017/04 Long night run 96121.66890
739811FE1_24CH_29650017/04 0.43 0.73 96921.56640
739823FE1_24CH_29730017/04 1.3 1.3 97021.97460
739827FE1_24CH_29720017/04 1.0 1.3 97021.97350
739830FE1_24CH_29610017/04 0.27 0.54 96921.86240
739833FE1_24CH_29700017/04 0.82 1.04 97021.77150
739965FE1_24CH_29710020/04 after adding br to the chamber 1.13 2.58 96723.37150

Session APRIL2: 3/4

We put in 3/4 position. It s between the Low Radius one and central one to guarantee the full acceptance. All 4 Febs are on. We start on 17/04. In fat we realized later that there was a non efficient region at the end of th scintillator. So we repeated some measurement, see below.

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739844FE1_24CH_29730017/04 18h25 1.00 1.68 96921.67450
739845FE1_24CH_29720017/04 20h05 1.00 1.27 96921.27350
739846FE1_24CH_29710017/04 22h45 0.90 1.27 96921.27240
739847FE1_24CH_297000Ib, iT slowly increase during the night 0.85-1.4 1.27-1.4 96921.2 7140
739848FE1_24CH_29690018/04 08h32 0.82 1.37 973 21.8 7030
739850FE1_24CH_29670018/04 10h44 0.74 1.13 973 21.8 6830
739854FE1_24CH_29650018/04 11h43 0.66 0.96 974 21.8 6620

Session APRIL3: Noise

Now runs collected with radom trigger of few kHz for the noise:

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739861FE1_24CH_296100 97221.66220
739862FE1_24CH_296500 0.5 0.9 97221.66620
739863FE1_24CH_296700 0.65 1.0 97221.66830
739864FE1_24CH_296900 May be 6800? 0.8 1.2 97221.67030
739865FE1_24CH_296900 0.93 1.32 97321.47030
739866FE1_24CH_297000 1.04 1.55 97321.47130
739869FE1_24CH_297100 1.1 1.8 97321.47240
739873FE1_24CH_297300 1.25 2.1 97321.47450
739875FE1_24CH_297200 1.1 1.8 97321.47340
739876FE1_24CH_29500 97121.7600
739885FE1_24CH_290 97121.70
739896FE1_24CH_290 Crate off 97322.20
739900FE1_24CH_290 Crate off, cables disconnected 97322.20

Session APRIL4: 3/4 with full acceptance

We realized looking on GAP footprint (black line) that there might be some acceptance problems for edge of the scintillator. So we moved the scintillators backward and ~5 cm closer to the center to be sure that we have full acceptance. Restart few runs with cosmic trigger.

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739886FE1_24CH_297100 1.0 1.24 97021.47240
739888FE1_24CH_296900 0.7 1.15 97021.37040
739889FE1_24CH_29670021.45 0.54 1.22 97121.56830
739890FE1_24CH_297300 23.051.15 2.18 97221.77450
739892FE1_24CH_296500 9.57 19/04 0.52 1.34 973 22.06630
739906FE1_24CH_296800 11.47 19/04 0.52 1.34 972 21.86930

Session APRIL5: threshold scan

Change threshold to 510 to see how noise goes down.

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739910FE1_24CH_297100 Thr 510 97221.67240
739927FE1_24CH_297100 Thr 514 97221.67240
739932FE1_24CH_297100 Thr 514 Noise run 97021.87240
739935FE1_24CH_297100 Thr 510 Noise run 97021.87240

Session APRIL6: radioactive source scan

We use CS 137 radioactive point like source with an activity of few MBq

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739937FE1_24CH_297100 Thr 510 Noise run source 1/4 position 97021.87240
739939FE1_24CH_297100 Thr 510 Noise run source middle position 97021.87240

Session APRIL7: Noise

Th chamber was refurbshed wit bars to make it eligible for the trolley. First we do a noise scan at THR = 506.

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739954FE1_24CH_290 Crate off, cables disconnected 97023.50
739955FE1_24CH_290 96923.70
739958FE1_24CH_29500 96923.60
739959FE1_24CH_296500 0.55 1.15 97023.50
739960FE1_24CH_296700 0.661.14 96923.10
739961FE1_24CH_296900 0.981.40 96822.70
739962FE1_24CH_297100 1.31.64 96822.60
739963FE1_24CH_297200 1.262.4 96722.90

Session APRIL8: Eff test after adding bars

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739965FE1_24CH_29710020/04 after adding br to the chamber 1.13 2.58 96723.37150

Session APRIL9: T0 resolution scan

we use now KODEL scintillators in synch. positioned above the setup. Ths run is not good for efficiency only for T0 resolution. The rate is ~0.2 Hz. Position 1/4.

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739970FE1_24CH_297100 1.22 2.4 96723.8???

The run above had a fundamental hassle that 2 KODEL scintillators 30 cm above the chamber covered with muons a large part of the chamber that is bad for dT0 resolution. So after some attempts we added a large scintillator below the detector to reduce the angle of detected muons. The rate went down to 1 muon / minute.

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739972FE1_24CH_297100 Currents of bottom gap rising 1.52 3.66 96723.8???

Session APRIL10: NOISE THR = 491 + 15 = 506 units

run DAQ state HVapp comment IBot ITop Eff P T HVeff
739974FE1_24CH_297100 19.20-19.30 / 23.04 - leds 1.4 3.60 96523.3
739975FE1_24CH_297100 19.33-19.43 / 23.04 - no leds 1.6 3.50 96423.1
new_measurement_cern_0318.txt · Dernière modification: 2018/04/23 19:40 par DAQ user